OpenStudy (anonymous):

If f(x) = x2 and g(x) = x + 6, find g(f(0)).

OpenStudy (akashdeepdeb):

f(x) = x^2 thus, f(0) = (0)^2 = 0 [Keep this aside for a while] g(x) = x+6 g(f(0)) = g(0) [Because f(0) = 0] g(0) = (0) + 6 = 6 Did you UNDERSTAND this? :D

OpenStudy (anonymous):

No i do not understand this help me...

OpenStudy (anonymous):

(:

OpenStudy (akashdeepdeb):

Alright look! :D Let us take one example: f(x) = x + 100 Now what is f(200) and f(2) and f(y) ?

OpenStudy (akashdeepdeb):

@Marifer

OpenStudy (anonymous):

Im taking a pretest and need the answer,

OpenStudy (akashdeepdeb):

If YOU want the answer, you NEED to get the concept. I am ready to teach you now, are you ready to learn?

OpenStudy (isaiah.feynman):

@AkashdeepDeb seems there is no time to learn, at the this point is assumed you only know the concepts

OpenStudy (akashdeepdeb):

Well, I actually have already given the answer. :P